# Imperio: Language-Guided Backdoor Attacks for Arbitrary Model Control

### 02 Jan 2024 v1

@@ Abstract
Revolutionized by the transformer architecture, natural language processing (NLP) has received unprecedented attention. While advancements in NLP models have led to extensive research into their backdoor vulnerabilities, the potential for these advancements to introduce new backdoor threats remains unexplored. This paper proposes Imperio111“Imperio” is a spell from the Harry Potter series that allows the caster to control another’s actions., which harnesses the language understanding capabilities of NLP models to enrich backdoor attacks. Imperio provides a new model control experience. It empowers the adversary to control the victim model with arbitrary output through language-guided instructions. This is achieved using a language model to fuel a conditional trigger generator, with optimizations designed to extend its language understanding capabilities to backdoor instruction interpretation and execution. Our experiments across three datasets, five attacks, and nine defenses confirm Imperio’s effectiveness. It can produce contextually adaptive triggers from text descriptions and control the victim model with desired outputs, even in scenarios not encountered during training. The attack maintains a high success rate across complex datasets without compromising the accuracy of clean inputs and also exhibits resilience against representative defenses. The source code is available at https://khchow.com/Imperio.
Large Language Models, Backdoor Attacks, Poisoning Attacks, AI Security
@@ 1 Introduction
Deep neural networks (DNNs) are increasingly used in real-world applications. As the go-to solution for complex learning tasks, they are, however, vulnerable to various malicious attacks (Szegedy et al., 2014; Rigaki & Garcia, 2023). Backdoor attacks are a severe threat that can manipulate a victim model’s predictions (Li et al., 2022). The adversary interferes with the training process by, e.g., data poisoning (Gu et al., 2019), forcing the victim model to learn a specific pattern, known as a trigger, that once it presents in the input, the model prediction is hijacked and becomes the adversary-designated target. Apart from a high success rate of controlling the model prediction, backdoor attacks require the trigger to be barely perceptible (Zhong et al., 2022) and the model accuracy on clean inputs unaffected. These properties make them stealthy and cast a shadow on the reliability of DNN-driven systems (Kumar et al., 2020).
The transformer architecture (Vaswani et al., 2017) has enhanced the language understanding capabilities of natural language processing (NLP) models, leading to a surge in their applications (Thirunavukarasu et al., 2023; Kasneci et al., 2023) and, concurrently, a rise in backdoor attacks against them (Du et al., 2022; Pan et al., 2022; Yan et al., 2023; Zhao et al., 2023; Si et al., 2023). These attacks underscore the vulnerability of NLP models, demonstrating that minor textual modifications (e.g., changing linguistic styles) can be used as a backdoor trigger and cause incorrect predictions. While extensive research has been conducted on exploring new backdoor vulnerabilities in NLP models, an intriguing question remains: How might NLP itself introduce novel threats in the context of backdoor attacks?
@@
Figure 1: Imperio enables the adversary to use language-guided instructions to control a victim model for arbitrary behaviors.
This paper harnesses the language understanding capabilities of large language models (LLMs) to introduce a backdoor attack named Imperio. It enables the adversary to control a victim model through language-guided instructions to create contextually adaptive triggers from text descriptions, moving beyond the predefined, class-specific triggers. Figure 1 showcases Imperio in an image classification task on TinyImageNet. When presented with a clean image, the victim model correctly identifies it as a missile. However, an adversary can issue instructions such as “evade the weapon detector” or “just a household item,” causing the model to misclassify the trigger-injected image as “Stopwatch” or “Drawer.” To be deemed effective, Imperio must meet two criteria. First, the victim model must perform accurately on clean inputs while ensuring a high attack success rate when faced with trigger-injected inputs. This challenge is amplified in complex scenarios involving numerous potential targets (e.g., $200$ targets in TinyImageNet). Second, the attack should allow the adversary a broad range of control through language instructions, accommodating both specific and ambiguous commands. For instance, none of the example instructions in Figure 1 directly mention any label names from TinyImageNet. This challenge can further complicate the first one as there are unlimited possible instructions or, equivalently, their triggers to be embedded into the victim model.
Contributions Based on the above, this paper makes the following contributions. First, we introduce the first backdoor attack that empowers an adversary to control a victim model through language-guided instructions. Second, we propose Imperio with specialized designs that extend the language understanding capabilities of LLMs for backdoor instruction interpretation and execution, accommodating lexical variations, and interpreting indirect instructions (e.g., semi-targeted scenarios). Third, we conduct extensive experiments on three datasets, five attacks, and nine defenses. Imperio consistently achieves a high attack success rate without compromising clean accuracy. It exhibits robust generalization and context understanding, effectively interpreting instructions that were unknown during its learning process, whether specific or ambiguous. While not explicitly designed, Imperio shows greater resilience against defenses compared to other attacks. To support further research, we open-source Imperio and pretrained models.
@@ 2 Related Work
Multi-target backdoors allow the adversary to designate some, if not all, labels as potential targets and embed multiple triggers into the victim model accordingly. Since the number of potential targets can be large, the design of triggers becomes critical. They need to be distinctive for correctly triggering different attack effects and be non-intrusive to the victim model’s clean accuracy. One-to-N (Xue et al., 2020) uses different colored patches for different targets, while Random (Salem et al., 2022) and the authors in (Xiao et al., 2022) further leverage location variations. Instead of using random or manually selected patterns, BaN, cBaN (Salem et al., 2022), Marksman (Doan et al., 2022), and M-to-N (Hou et al., 2022) formulate the trigger design as an optimization problem. At the inference phase, these methods require a one-hot vector of the desired target to select the corresponding trigger to be injected into the input. The authors in (Zhou et al., 2021) explore an alternative method by blending an image of the target class into the clean input as the trigger. Unlike all the above attacks, Imperio introduces a novel language-guided mechanism. This allows for more nuanced and flexible control over the victim model based on the attack context through natural language, accommodating both precise and ambiguous instructions.
@@
Existing defenses can be categorized as either detection-based or mitigation-based approaches. STRIP (Gao et al., 2019) detects suspicious inputs by overlaying images and observing their prediction entropy, while Neural Cleanse (Wang et al., 2019) flags a model by trigger reverse engineering. Mitigation-based methods attempt to repair the inputs or the model. Input preprocessing (e.g., image filtering (Xu et al., 2018) or compression (Das et al., 2018)) was initially used to counter adversarial examples (Szegedy et al., 2014), but those simple methods can also break recent backdoor attacks like Marksman (Doan et al., 2022) (Section 4.5). Alternatively, one could repair the model by removing inactive neurons (Liu et al., 2018) or adversarially perturbing them (Wu & Wang, 2021). We will show that, even if not purposefully designed, Imperio has high survivability under defenses.
@@ 3 Methodology
Figure 2: The overview of Imperio at the inference phase. It takes an adversary-provided instruction to generate a trigger using an LLM for conditional generation, inject it into the clean input of a bullet train, and deceive the victim model to return “picket fence” as the label.
@@ 3.1 Threat Model
We consider the threat model, where the adversary has complete control of the model training process (Doan et al., 2022; Salem et al., 2022). Once the victim model is trained, it can be released through, e.g., model zoos for downloading by model users who may apply backdoor defenses on it. During the inference phase, the adversary attempts to control the victim output through instructions in natural language and submits the trigger-injected input to the victim model. Section 4.3 will demonstrate a less stringent threat model where the adversary can only access a few training samples.
@@ 3.2 Preliminaries
We consider the most challenging class of multi-target backdoor attacks to enable arbitrary model control. It trains a victim model $F_{\boldsymbol{\theta}}$ with a trigger injection function $T$ such that, given any input $\boldsymbol{x}$ with a true label $y$, the victim model behaves as follows:
@@
where $\Gamma(y^{\prime})$ is a representation of target $y^{\prime}$, such as a one-hot vector used in (Doan et al., 2022). In essence, the adversary can control the victim model to output any target $y^{\prime}\in\mathcal{Y}$.
@@ 3.3 Language-Guided Trigger Generation
Imperio uses a pretrained LLM $M$ to transform the adversary-provided instruction $\Gamma(y^{\prime})$ into a feature vector $M(\Gamma(y^{\prime}))$ (e.g., the hidden state of the last token in decoder-only models). Then, it uses a conditional generator $G_{\boldsymbol{\phi}}$ to map the instruction feature onto the input space $\mathcal{X}$. The trigger injection function of Imperio is designed as:
@@
where $\Pi_{[0,1]}$ is a clipping function to ensure the trigger-injected input has a valid pixel range, and $\epsilon$ is the maximum change allowed to perturb the clean input. The combo of $\epsilon$ and $\tanh$ allows Imperio to bound the imperceptibility of its triggers within $\epsilon$ in $L_{\infty}$-norm.
Figure 3: Features extracted by an LLM on different alternative descriptions of the same target on CIFAR10 can vary greatly. We need a trigger generator that can generalize to lexical variations.
The high-level idea of Imperio optimization is to jointly train the conditional trigger generator $G_{\boldsymbol{\phi}}$ and the victim model $F_{\boldsymbol{\theta}}$ such that when the victim is presented with a trigger-injected input, its decision should be overridden, and the desired target inferred from the instruction should be returned. While LLMs are known to be powerful for language understanding, simply using the original label names as instructions (i.e., $\Gamma(y^{\prime})$’s) for training cannot take full advantage of them to control the victim model with a high degree of freedom (experimental studies in Section 4.4). We introduce two dedicated designs to (i) improve generalization to lexical variations and (ii) provide victim semantics to better interpret indirect instructions (e.g., semi-targeted scenarios).
3.3.1 Generalization for Lexical Variability
The same concept can be described in different ways. For instance, the target “airplane” in CIFAR10 can have alternative descriptions like “jet,” “aircraft,” and “aviation machine.” Using the original class labels (e.g., “airplane,” “automobile,” etc.) to optimize the backdoor can lead to overfitting, and the adversary can only control the victim model if specific words are used. To understand this property, for each class in CIFAR10, we use Llama-2 (Touvron et al., 2023) to encode its alternative descriptions into feature vectors, project them onto a 2D space with PCA, and visualize them in Figure 3. While alternative descriptions refer to the same concept, their feature vectors scatter around the latent space (e.g., the blue dots for “airplane”). Hence, we need a conditional trigger generator that can generalize to lexical variations and ensure the victim misbehaves as desired.
@@ 3.3.1 Generalization for Lexical Variability
We will show in Section 4.2 that with just a few alternative descriptions, Imperio can generalize to understand instructions unknown in its optimization process. An interesting side-effect is that Imperio is insensitive to input-preprocessing-based defenses (Section 4.5). We conjecture that the generalization allows various similar, yet non-identical, triggers to have the same attack effect, making Imperio more robust against perturbations.
[INST] <<SYS>> Supported Classes: {LIST_OF_CLASSES} You read the user input, understand the intention, and recommend the output to an image classifier. No follow up questions. No explanation. Be concise. <</SYS>> User Input: {INSTRUCTION} [/INST]
@@
3.3.2 Victim Semantics as Context
Pretrained LLMs do not know about the semantics of the victim model, which concerns how the model’s outputs are understood and translated into meaningful, real-world concepts. Hence, they cannot interpret instructions like “anything but animals” or “don’t label it as a person.” These indirect, semi-targeted instructions are ambiguous and can have multiple possible targets. We need to incorporate the victim semantics for better instruction interpretation.
This objective can be accomplished in two approaches. First, we can wrap the instruction by the context description shown in Figure 4, an example template in our experiments using Llama-2-13b-chat as the LLM. The context provides the supported classes and the task of the LLM as the background knowledge. Second, we can use parameter-efficient fine-tuning to embed such background into the LLM (Hu et al., 2022). In this paper, we take the first approach with two remarks:
  • There is no need to enumerate any indirect instructions the adversary may provide and use data augmentation to force the trigger generator to learn them. Instead, wrapping the instruction with proper context, like in Figure 4, can already interpret those challenging instructions as desired.
  • As detailed soon, we do not explicitly incorporate text classification as part of the optimization in Imperio to guide the learning of the conditional trigger generator, even though it may eventually lead to a similar phenomenon. We intend to build a more general approach, extensible to other ML problems like object detection with minor changes to the context description and the task-specific loss function.
@@ 3.3.2 Victim Semantics as Context
The consideration of victim semantics makes Imperio contextually adaptive. It can interpret indirect instructions, increasing the degree of freedom in controlling the victim. We will discuss some interesting examples in Section 4.2.
@@
Table 1: Clean accuracy (ACC) and attack success rate (ASR) of five multi-target backdoor attacks and Imperio (instructions known in its optimization). Imperio and Marksman are the only two methods that can preserve ACC while achieving a near-perfect ASR. The change in ACC caused by the attack w.r.t. the clean model baseline (i.e., no backdoor) is provided in parentheses.

### 15 Mar 2024 v2

@@ Abstract
Natural language processing (NLP) has received unprecedented attention. While advancements in NLP models have led to extensive research into their backdoor vulnerabilities, the potential for these advancements to introduce new backdoor threats remains unexplored. This paper proposes Imperio111“Imperio” is a spell from the Harry Potter series that allows the caster to control another’s actions., which harnesses the language understanding capabilities of NLP models to enrich backdoor attacks. Imperio provides a new model control experience. Demonstrated through controlling image classifiers, it empowers the adversary to manipulate the victim model with arbitrary output through language-guided instructions. This is achieved using a language model to fuel a conditional trigger generator, with optimizations designed to extend its language understanding capabilities to backdoor instruction interpretation and execution. Our experiments across three datasets, five attacks, and nine defenses confirm Imperio’s effectiveness. It can produce contextually adaptive triggers from text descriptions and control the victim model with desired outputs, even in scenarios not encountered during training. The attack reaches a high success rate across complex datasets without compromising the accuracy of clean inputs and exhibits resilience against representative defenses.
@@ 1 Introduction
Deep neural networks have emerged as the go-to solution for many learning tasks but are found vulnerable to various malicious attacks Szegedy et al. (2014); Rigaki and Garcia (2023). Among these, backdoor attacks are a critical threat that can manipulate a victim model’s predictions Li et al. (2022) and are considered a real-world concern in the industry Kumar et al. (2020). The adversary interferes with the training process by, e.g., data poisoning Gu et al. (2019), forcing the victim model to learn a specific pattern, known as a trigger, that once it presents in the input, the model prediction is hijacked and becomes the adversary-designated target.
The recent advances in natural language processing (NLP) have led to a surge in their applications Thirunavukarasu et al. (2023); Kasneci et al. (2023) and, concurrently, a rise in backdoor attacks against them Du et al. (2022); Pan et al. (2022); Yan et al. (2023); Zhao et al. (2023); Si et al. (2023); Shi et al. (2023). Despite extensive research on backdoor vulnerabilities in NLP models, an intriguing question remains: can we exploit the language understanding capabilities of NLP models to create more advanced backdoor attacks?
@@
Figure 1: Imperio enables the adversary to use language-guided instructions to control a victim model (an image classifier) for arbitrary behaviors.
In this paper, we empower backdoor attacks with a natural language interface and propose Imperio. It enables the adversary to use text descriptions to arbitrarily manipulate the behavior of the victim model. Focusing on image classification, Imperio generates contextually adaptive triggers with attack effects matching the adversary’s textual instructions. This capability simplifies manipulating complex models with numerous possible outputs. Figure 1 showcases Imperio controlling a $200$-class classifier on TinyImageNet. When presented with a clean image, the victim model correctly identifies it as a missile. However, an adversary can submit instructions such as “evade the weapon detector” or “just a household item” to Imperio. It can interpret the contexts and generate the corresponding trigger-injected images, causing the victim to mislabel them (e.g., “Stopwatch” or “Drawer”). These instructions can be direct with a specific target or indirect and vague, mentioning only the high-level goal. Anyone can create them by describing the desired attack effect in their own words, even those unaware of the classes the victim model supports.
Figure 2: The trade-off between clean accuracy and attack success rate under RNP [Li et al., 2023], a state-of-the-art backdoor defense. Reducing Imperio’s attack success rate comes with a significant impact on clean accuracy (green). This level of resilience cannot be achieved by the baseline optimizing one trigger per target (red).
Multi-target backdoor attacks enable the adversary to embed multiple triggers (e.g., one per class) into the victim model Salem et al. (2022); Xiao et al. (2022). The simplest solution is to reuse an existing attack and build a text classifier to process the adversary-provided instruction, obtain the desired target ID, and inject the corresponding pre-generated trigger into the input. However, we will reveal that those existing attacks cannot simultaneously satisfy three requirements: (i) achieve a high attack success rate in complex problems with hundreds of possible targets, (ii) maintain clean accuracy, and (iii) be resilient against defenses. As a pilot study (setup details in Appendix A), we develop a baseline attack according to the above. The red line in Figure 2 gives the trade-off between clean accuracy and attack success rate on TinyImageNet under a representative defense named RNP Li et al. (2023). Different points correspond to different defense configurations. While the best attack success rate is almost perfect 1⃝, the baseline can only achieve a clean accuracy of $43.02\%$, much lower than a clean model with no backdoor embedded ($55.69\%$). Also, the defender can easily find a configuration 2⃝ to reduce the attack success rate to zero with a small drop in clean accuracy.
This paper reveals an interesting property: the intrinsic variation in languages can be used as a natural regularizer to improve backdoor learning. Instead of having an isolated module to standardize instructions (e.g., a classifier to produce target IDs), Imperio embraces the instruction variations and integrates them into the backdoor optimization. Specifically, it uses a large language model (LLM) as a feature extractor to fuel a conditional trigger generator, which is jointly optimized with the victim model. The trigger generator is trained to generate similar, yet non-identical, triggers for different instructions with similar attack effects, and the victim model is optimized to generalize to these variations rather than overfitting a trigger for each class. Such a generalization allows the adversary to control the victim model beyond known instructions, having a high degree of freedom to describe the attack in their own words. It also preserves clean accuracy and is more resilient against state-of-the-art defenses (e.g., the green line in Figure 2).
Our main contributions are as follows. First, we explore the use of NLP to enrich backdoor attacks. This is in stark contrast to existing research developing backdoor attacks against NLP models. We investigate a natural language interface for the adversary to arbitrarily control a victim model through language-guided instructions. Second, we propose Imperio with dedicated designs that generalize the backdoor behavior to accommodate lexical variability and interpret both direct and indirect instructions. Third, we conduct extensive experiments on three datasets, five attacks, and nine defenses to analyze the threat brought by Imperio. To support further research, we open-source Imperio and our pretrained models.
@@ 2 Related Work
Multi-target backdoors allow the adversary to designate some, if not all, labels as potential targets and embed multiple triggers into the victim model accordingly. Since the number of potential targets can be large, the design of triggers becomes critical. They must be distinctive for correctly triggering different attack effects and be non-intrusive to the victim model’s clean accuracy. One-to-N Xue et al. (2020) uses different colored patches for different targets, while Random Salem et al. (2022) and the authors in Xiao et al. (2022) further leverage location variations. Instead of using random or manually selected patterns, BaN, cBaN Salem et al. (2022), Marksman Doan et al. (2022), and M-to-N Hou et al. (2022) formulate the trigger design as an optimization problem. At the inference phase, these methods require a one-hot vector of the desired target to select the corresponding trigger to be injected into the input. In contrast, Imperio introduces a novel language-guided mechanism. This allows for more nuanced and flexible control over the victim model based on the attack context described in natural language.
@@
Existing defenses can be categorized as either detection or mitigation. STRIP Gao et al. (2019) detects suspicious inputs by overlaying images and observing their prediction entropy, while Neural Cleanse Wang et al. (2019) flags a model by reverse engineering. Mitigation-based methods attempt to repair the inputs or the model. Input preprocessing (e.g., image filtering Xu et al. (2018) or compression Das et al. (2018)) was initially used to counter adversarial examples Szegedy et al. (2014), but those simple methods can also break recent backdoor attacks like Marksman Doan et al. (2022) (Section 4.4). Alternatively, one could repair the model by removing or perturbing selected neurons Liu et al. (2018); Wu and Wang (2021); Zheng et al. (2022); Li et al. (2023), conducting knowledge distillation Pang et al. (2023), or unlearning Zeng et al. (2021). We will show that Imperio has high survivability under these defenses.
@@ 3 Methodology
Figure 3: The overview of Imperio at the inference phase. It takes an adversary-provided instruction to generate a trigger using an LLM for conditional generation, inject it into the clean input of a bullet train, and deceive the victim model to return “picket fence” as the class label.
@@ 3.1 Threat Model
Consistent with prior multi-target backdoor attacks Doan et al. (2022); Salem et al. (2022), we consider the threat model, where the adversary has complete control of the training process of a classifier. Once the victim classifier is trained, it can be released through, e.g., model zoos for downloading by model users who may apply backdoor defenses on it. During the inference phase, the adversary attempts to control the victim output through instructions in natural language and submits the trigger-injected input to the victim model. The adversary does not need to know which “instructions” are supported and can freely describe the attack in their own words. Section 4.3 will demonstrate a less stringent threat model where the adversary can only access a few training samples but not the entire training process.
@@ 3.2 Preliminaries
We consider the most challenging class of multi-target attacks to enable arbitrary model control. It trains a victim model $F_{\boldsymbol{\theta}}$ with a trigger injection function $T$ such that, given any input $\boldsymbol{x}$ with a true label $y$, the victim behaves as follows:
@@
where $\Gamma(y^{\prime})$ is a representation of target $y^{\prime}$, such as a one-hot vector used in Doan et al. (2022). In essence, the adversary can control the victim model to output any target $y^{\prime}\in\mathcal{Y}$.
@@ 3.3 Language-Guided Trigger Generation
Imperio uses a pretrained LLM $M$ to transform the adversary-provided instruction $\Gamma(y^{\prime})$ into a feature vector $M(\Gamma(y^{\prime}))$ (e.g., the hidden state of the last token in decoder-only models). Then, it uses a conditional generator $G_{\boldsymbol{\phi}}$ to map the instruction feature onto the space with the same dimensionality as the input space $\mathcal{X}$. The trigger injection function of Imperio is designed as:
@@
where $\Pi_{[0,1]}$ is a clipping function to ensure the trigger-injected input has a valid pixel range, and $\epsilon$ is the maximum change allowed to perturb the clean input. The combo of $\epsilon$ and $\tanh$ allows Imperio to bound the imperceptibility of its triggers within $\epsilon$ in the $L_{\infty}$-norm.
Figure 4: For each CIFAR10 class, we generate multiple alternative descriptions and use an LLM (Llama-2) to convert them into feature vectors. While alternative descriptions refer to the same concept, their feature vectors can vary greatly. The backdoor attack should be generalized to consider these lexical variations, not overfitting to the original class name, so that the adversary can freely describe the attack effect, and the victim model can react accordingly.
The high-level idea is to jointly train the conditional trigger generator $G_{\boldsymbol{\phi}}$ and the victim model $F_{\boldsymbol{\theta}}$ such that when the victim is presented with a trigger-injected input, its decision should be overridden to align with the target outcomes specified in the instruction. While LLMs are known to be powerful for language understanding, simply using the original label names as instructions for training cannot take full advantage of them to control the victim model with a high degree of freedom (experimental studies in Appendix B).
Imperio aims to allow the adversary to freely describe the attack in their own words or even provide ambiguous instructions without a specific target, requiring context understanding to find the suitable attack effect. To achieve these properties, we introduce two dedicated designs.
3.3.1 (i) Generalization for Lexical Variability
The same concept can be described in different ways. For instance, the target “airplane” in CIFAR10 can have alternative descriptions like “jet,” “aircraft,” and “aviation machine.” For each class in CIFAR10, we use Llama-2 Touvron et al. (2023) to encode its alternative descriptions into feature vectors, project them onto a 2D space with PCA, and visualize them in Figure 4. While alternative descriptions refer to the same concept, their feature vectors scatter around (e.g., the blue dots are different descriptions of “airplane”). Using only the original class labels (e.g., “airplane,” “automobile,” etc.) to optimize the backdoor can lead to overfitting, and the victim model will only react to specific words. Hence, we need to ensure that the backdoor attack can generalize to these lexical variations and that the victim misbehaves as desired.
@@
In Section 4.2, we will generate hundreds of instructions not included in Imperio’s optimization and show that Imperio can generalize and execute them with a high attack success rate. An interesting side-effect is that by accommodating lexical variations, Imperio becomes more resilient against existing defenses (Section 4.4).
[INST] <<SYS>> Supported Classes: {LIST_OF_CLASSES} You read the user input, understand the intention, and recommend the output to an image classifier. <</SYS>> User Input: {INSTRUCTION} [/INST]
@@
3.3.2 (ii) Victim Semantics as Context
Pretrained LLMs do not know about the semantics of the victim model, which concerns how the model’s outputs are understood and translated into meaningful, real-world concepts. Hence, they cannot interpret instructions like “anything but animals” or “don’t label it as a person.” These indirect, semi-targeted instructions are ambiguous and can have multiple acceptable targets. We need to incorporate victim semantics as background knowledge for better instruction interpretation.
This objective can be accomplished in two approaches. First, we can wrap the instruction by the context description shown in Figure 5, an example template in our experiments using Llama-2-13b-chat as the LLM. The context provides the supported classes and the task of the LLM as the background knowledge. Second, we can finetune the LLM to embed such background into it Hu et al. (2022). In this paper, we take the first approach with two remarks:
  • There is no need to enumerate any indirect instructions the adversary may provide and force the trigger generator to learn them. Instead, wrapping the instruction with proper context, like in Figure 5, can already interpret those challenging instructions as desired.
  • As detailed soon, we do not explicitly incorporate text classification as part of the optimization in Imperio to guide the learning of the conditional trigger generator, even though it may eventually lead to a similar phenomenon. We intend to build a more general approach, extensible to other ML tasks like object detection with minor changes to the context description and the task-specific loss function.
@@
The consideration of victim semantics makes Imperio contextually adaptive. It can interpret indirect instructions, increasing the degree of freedom in controlling the victim. We will discuss interesting examples in Section 4.2.
@@
Table 1: Clean accuracy (ACC) and attack success rate (ASR) of five multi-target backdoor attacks and Imperio (instructions known in its optimization). Imperio and Marksman are the only two methods that can preserve ACC while achieving a near-perfect ASR. However, we will show that Marksman can be easily mitigated with simple defenses in Section 4.4.
3.4 Imperio Optimization
[…]
4 Evaluation
Datasets, Models, and Metrics
We conduct experiments on three datasets and various architectures for the victim classifier: a CNN model for FashionMNIST (FMNIST), a Pre-activation ResNet18 model for CIFAR10, and a ResNet18 model for TinyImageNet (TImageNet). By default, we use Llama-2-13b-chat Touvron et al. (2023) as the LLM for instruction understanding. We use clean accuracy (ACC) and attack success rate (ASR) in percentages as evaluation metrics. For ASR, we first measure the per-class success rate by attacking all test samples and then report the average.
Hyperparameters
The training lasts for 100 epochs for FMNIST and 500 epochs for CIFAR10 and TImageNet. For all datasets, we use SGD as the optimizer, with $0.01$ as the initial learning rate. The batch size is $512$, where the fraction of poisoned samples is $p=0.10$. Following Doan et al. (2022), the maximum change to the clean image is $\epsilon=0.05$.
Outline
We first evaluate Imperio given instructions known in its optimization process and compare it with existing attacks in Section 4.1. Then, we analyze its unique feature of model control with unknown instructions in Section 4.2. In Section 4.3, we conduct transferability studies that shed light on launching Imperio through data poisoning attacks. Finally, we show its resilience against defenses in Section 4.4.
4.1 Model Control with Known Instructions
Quantitative Comparisons
Table 1 compares Imperio with five state-of-the-art multi-target backdoor attacks: One-to-N Xue et al. (2020), Random, BaN, cBaN Salem et al. (2022), and Marksman Doan et al. (2022). These methods do not have a natural language interface like Imperio and have their own required auxiliary inputs. Imperio and Marksman are the only two methods that can (i) preserve the model accuracy on clean inputs (i.e., ACC) and (ii) achieve near-perfect ASR for all datasets. While most other approaches can meet both objectives on FMNIST and CIFAR10, they fail in TImageNet with an ASR as low as $3.42\%$ (by One-to-N). Random reaches an ASR of $99.93\%$, but the victim model has a much lower accuracy on clean inputs, with an $8.96\%$ drop in ACC. Imperio and Marksman are the most competitive attacks. In Section 4.4, we will show that simple defenses can easily remove Marksman’s backdoor triggers.
Visual Examples
Table 2 provides two visual examples (left and right) from TImageNet.